OpenStudy (anonymous):

How to graph this system of equations? x + 2y = –10 3x + y = 12

OpenStudy (anonymous):

find the x and why intercepts, and connect them. To find the X-axis intercept, change the y-value to 0. to find the Y-axis intercept, change the x-value to 0. do that for both lines.

OpenStudy (anonymous):

can you show please? (:

OpenStudy (anonymous):

\[x+2(0) = -10; \therefore x = -10, thus (-10,0) \] (-10,0) would be a point on that line. \[(0)+2y=-10; y = -10/2; y = -5\] therefore the point (0,-5) would be on the line. connect the two points and you have a line.

OpenStudy (anonymous):

do the same fore the other line.